### v2

# Investigating Popularity Bias Amplification in Recommender Systems Employed in the Entertainment Domain

Abstract.
Recommender systems have become an integral part of our daily online experience by analyzing past user behavior to suggest relevant content in entertainment domains such as music, movies, and books. Today, they are among the most widely used applications of AI and machine learning. Consequently, regulations and guidelines for trustworthy AI, such as the European AI Act, which addresses issues like bias and fairness, are highly relevant to the design, development, and evaluation of recommender systems. One particularly important type of bias in this context is popularity bias, which results in the unfair underrepresentation of less popular content in recommendation lists. This work summarizes our research on investigating the amplification of popularity bias in recommender systems within the entertainment sector. Analyzing datasets from three entertainment domains, music, movies, and anime, we demonstrate that an item’s recommendation frequency is positively correlated with its popularity. As a result, user groups with little interest in popular content receive less accurate recommendations compared to those who prefer widely popular items. Furthermore, this work contributes to a better understanding of the connection between recommendation accuracy, calibration quality of algorithms, and popularity bias amplification.
Recommender Systems, Fairness, Popularity Bias, Bias Amplification, Collaborative Filtering
copyright: ccjournalyear: 2025conference: Fourth European Workshop on Algorithmic Fairness; June 30–July 02, 2025; Eindhoven, NL
1. Introduction and Motivation
Recommender systems have become one of the most prevalent applications of machine learning and AI, shaping our daily online experiences. They play a crucial role in helping users navigate vast and complex information spaces by identifying relevant content (Burke et al., 2011; Jannach et al., 2016). Since their early implementations (Resnick et al., 1994), these systems have relied on analyzing past user behavior to construct user models and provide recommendations, not only for items like movies, music, and books, but also for social connections in online networks (Kowald et al., 2013; Lacic et al., 2015; Eirinaki et al., 2018).
A variety of techniques have been used to build these user models, including traditional methods such as collaborative filtering (CF) (Ekstrand et al., 2011), content-based filtering (Lops et al., 2010), and hybrid approaches (Burke, 2002), as well as more recent techniques based on latent representations (embeddings) and deep learning (Chen et al., 2023b). The entertainment domain is one of the key areas where recommender systems are widely deployed, assisting users in discovering movies, music, books, and other media. Given the increasing adoption of recommender systems in both research and industry (Jannach and Zanker, 2022), and their inherently human-centric nature, it is essential to consider existing regulations and requirements for trustworthy AI (Di Noia et al., 2022). Various institutions, including the European Commission, have defined trustworthiness through multiple dimensions, leading to regulatory frameworks such as the EU Artificial Intelligence Act(Commission, 2021), which emphasizes issues like bias and fairness in AI. These concerns are particularly relevant in the context of recommender systems, as highlighted by recent studies on trustworthy recommendation models(Fan et al., 2023). While bias and fairness in AI and machine learning have received significant research attention in recent years (Mehrabi et al., 2021; Scher et al., 2023), the replication and amplification of biases remain open challenges, especially in interactive systems (Friedman and Nissenbaum, 1996) and recommender systems in particular (Chen et al., 2023a). One of the most prevalent biases in CF-based recommender systems is popularity bias, which results in the underrepresentation of less popular content in personalized recommendations (Elahi et al., 2021; Abdollahpouri et al., 2021).
The present work summarizes our research on investigating the amplification of popularity bias in recommender systems within the entertainment domain. Analyzing datasets from three sectors, music, movies, and animes, we demonstrate that popularity bias disproportionately impacts user groups with little interest in popular content, as they receive less accurate recommendations compared to those who favor popular items (Kowald et al., 2020, 2021; Kowald and Lacic, 2022; Kowald et al., 2023). Furthermore, this work provides deeper insights into the relationship between recommendation accuracy, algorithmic calibration quality, and the amplification of popularity bias. Finally, we propose directions for future research.
2. Background and Methods
Research has demonstrated that recommendation algorithms, particularly those based on collaborative filtering (CF), exhibit a strong bias toward popular items, resulting in their overrepresentation in recommendation lists (Ekstrand et al., 2018; Elahi et al., 2021). Conversely, this leads to the unfair underrepresentation of less popular, long-tail items (Brynjolfsson et al., 2006; Park and Tuzhilin, 2008). Various metrics have been proposed in the literature to assess and analyze popularity bias from both item and user perspectives (Ahanger et al., 2022; Klimashevskaia et al., 2023). In the following, we describe some of these metrics used in our research, and give an overview of datasets for recommender systems employed in the entertainment domain.
2.1. Metrics to Investigate Popularity Bias Amplification
Our research focuses on three key methods for measuring popularity bias across user groups: (i) differences in recommendation accuracy, (ii) miscalibration, and (iii) popularity lift. The first approach involves a straightforward comparison of average recommendation accuracy between groups using the mean average error (MAE). In contrast, miscalibration and popularity lift require more complex calculations.
In general, calibration measures the alignment between a user profile $p$ and a corresponding recommendation list $q$ in terms of genre distribution (Steck, 2018). For example, if a user historically consumes 80% rock and 20% pop music, a calibrated recommendation list should reflect a similar distribution. While not explicitly a popularity bias metric, calibration is frequently used to assess and interpret popularity bias in recommendations (Abdollahpouri et al., 2019a, 2020). The concept of miscalibration represents the deviation between $p$ and $q$ (Lin et al., 2020), quantified using the Kullback-Leibler (KL) divergence between the genre distributions in $p$, i.e., $p(c|u)$, and in $q$, i.e., $q(c|u)$:
(1) $KL(p||q)=\sum_{c\in C}p(c|u)\log\frac{p(c|u)}{q(c|u)}$
where $C$ is the set of all genres in the dataset. A value of $KL(p||q)=0$ indicates perfect calibration, while higher values (approaching 1) signify increasingly miscalibrated recommendations. These values can be averaged for a given user group $g$. We term miscalibration as MC in this paper.
In contrast, popularity lift quantifies the extent to which recommendation algorithms amplify the popularity bias inherent in user profiles (Abdollahpouri et al., 2019b, 2020). Specifically, it measures the disproportionate recommendation of popular items to a given user group $g$. The metric is based on the group average popularity $GAP_{p}(g)$, which represents the mean popularity of items in the user profiles $p$ of group $g$. Similarly, $GAP_{q}(g)$ denotes the average popularity of recommended items for users in $g$. Popularity lift is then defined as:
(2) $PL(g)=\frac{GAP_{q}(g)-GAP_{p}(g)}{GAP_{p}(g)}$
A value of $PL(g)>0$ indicates that recommendations for group $g$ are skewed toward more popular items, whereas $PL(g)<0$ suggests an overrepresentation of less popular content. The ideal scenario is $PL(g)=0$, where the popularity distribution remains unchanged. We term popularity lift as PL in this paper.
2.2. Datasets and Recommender System Algorithms Employed in the Entertainment Domain
In our research, we analyze three datasets from the entertainment section, namely Last.fm representing the music domain, MovieLens representing the movie domain, and MyAnimeList, representing the anime domain. These datasets are described in more detail in our previous work (Kowald and Lacic, 2022; Kowald et al., 2023), are summarized in Table 1, and are freely available via Zenodo [^1] https://zenodo.org/records/7428435. As described in (Kowald and Lacic, 2022; Kowald et al., 2023), we split the users in every dataset in three equally-sized groups (1,000 users) based on their inclination towards popularity. We term the groups LowPop, MedPop, and HighPop.
Table 1. Statistics of the datasets, including the number of users ($|U|$), items ($|I|$), ratings ($|R|$), and distinct genres ($|C|$), as well as sparsity, average interactions per user/item, and the rating range ($R$-range) (Kowald and Lacic, 2022; Kowald et al., 2023).
Dataset $|U|$ $|I|$ $|R|$ $|C|$ $|R|/|U|$ $|R|/|I|$ Sparsity $R$-range
Last.fm 3,000 131,188 1,417,791 20 473 11 0.996 [1–1,000]
MovieLens 3,000 3,667 675,610 18 225 184 0.938 [1–5]
MyAnimeList 3,000 9,450 649,814 44 216 69 0.977 [1–10]
We analyze two well-known personalized recommendation algorithms used in the entertainment section, namely user-based, k-nearest-neighbor CF (UserKNN) (herlocker1999algorithmic]) and non-negative matrix factorization (NMF) (Luo et al., 2014). For the sake of reproducibility (Semmelrock et al., 2023, 2024), the implementation details of these algorithms and the complete source code to reproduce all of our research results are available via GitHub [^2] https://github.com/domkowald/FairRecSys.
3. Results and Findings
In Figure 1b, we show the correlation between the popularity of music artists and their recommendation frequency in our Last.fm dataset. We see that both algorithms tend to favor popular music artists in their recommendation lists (Kowald et al., 2020, 2021). For the sake of space, we omit the results for MovieLens and MyAnimeList, but similar results can be obtained for these domains as well (Kowald and Lacic, 2022; Kowald et al., 2023). This means that the higher the popularity of an item, the higher is also the probability that this item is recommended. Next, we investigate if this item popularity bias also negatively influences user groups with little interest into popularity (i.e., our LowPop group).
(a) UserKNN
(b) NMF
Figure 1. Correlation of music artist popularity and recommendation frequency in the Last.fm dataset. Both algorithms investigated tend to favor popular music artists (Kowald et al., 2020, 2021). Similar results can be obtained for the movie and anime domains (Kowald and Lacic, 2022; Kowald et al., 2023)
Next, Table 2 summarizes our results for the three entertainment datasets across the three user groups and the two algorithms. We see that the LowPop user group always receives the statistically significant (according to a t-test with $p<0.05$) worst accuracy results (MAE). In addition, this user group also receives the most miscalibrated (MC) and popularity biased (PL) results, which helps to better understand the connection between accuracy and these two metrics (Kowald and Lacic, 2022; Kowald et al., 2023). Only in the Last.fm dataset, the other user groups reach worse PL estimates, which we attribute to the special characteristics of this dataset, namely the high number of items (see Table1 and also (Kowald et al., 2020)).
Table 2. MAE, MC, and PL results for the LowPop, MedPop, and HighPop user groups. The worst (i.e., highest) results are highlighted in bold. Statistical significance (t-test between LowPop vs. MedPop and HighPop) is indicated by * for $p<0.05$.
Data Last.fm MovieLens MyAnimeList
Algorithm Metric MAE MC PL MAE MC PL MAE MC PL
UserKNN LowPop 54.32* 0.51* 0.52 0.80* 0.75* 0.64* 1.37* 0.92* 0.74*
MedPop 46.76 0.50 0.82 0.75 0.69 0.37 1.34 0.72 0.22
HighPop 49.75 0.45 0.80 0.72 0.62 0.20 1.31 0.63 0.08
NMF LowPop 42.47* 0.54* 0.10 0.75* 0.78* 0.57* 1.01* 0.91* 0.87*
MedPop 34.03 0.52 0.17 0.72 0.71 0.37 0.97 0.72 0.35
HighPop 41.14 0.48 0.33 0.70 0.63 0.22 0.95 0.63 0.13
4. Conclusion and Future Research Directions
In this work, we summarize our research on investigating the amplification of popularity bias in recommender systems within the entertainment sector. Analyzing datasets from music, movies, and anime, we show that recommendation frequency increases with item popularity, disadvantaging users with little interest in popular content. Additionally, we explore the interplay between recommendation accuracy, algorithmic calibration, and the amplification of popularity bias. Future research should investigate additional domains with respect to popularity bias amplification (e.g., news (Lacic et al., 2022)), and propose robust methods for mitigating popularity bias (e.g., (Klimashevskaia et al., 2022)).
Acknowledgements. This work was supported by the FFG COMET program. A full version of this article was accepted as a habilitation at Graz University of Technology in June 2024, and is available via Arxiv (Kowald, 2024).